# Path Pooling: Train-Free Structure Enhancement for Efficient Knowledge Graph Retrieval-Augmented Generation

### v1

@@ A.3.1 Different Path Search Algorithms
Figure 11 summarizes Hit@1 and F1-Score of positional reranking mechanism with different path search algorithms on Qwen2.5-7b. Consistent with previous results, all three path search algorithms are effective in enhancing the performance of SubgraphRAG. As the number of triples increases, the Hit@1 and F1-Score of SubgraphRAG show a initially rising and then declining trend. However, when enhanced with a positional reranking mechanism based on BFS path search algorithm, both metrics demonstrate a consistently upward trend. For instance, with 100 triples, the Hit@1 of SubgraphRAG is 45.77, while it is 47.52 with BFS, which brings about an increase of 1.75%. The result emphasizes the effectiveness of incorporating the structure information of KGs.
@@ A.3.2 Different Pooling Strategies
As illustrated in Figure 12, positional reranking mechanism with different pooling strategies all enhance SubgraphRAG on Qwen2.5-7b. Similar benefits are observed when leveraging triple reselection mechanism and on Llama3.1-8b. As commonly used pooling strategies, both max pooling and average pooling demonstrate significant roles here. However, surprisingly, when the number of triples increases from 50 to 200, the positional reranking mechanism using max pooling achieves better enhancement effect than using average pooling. For example, with 50 triples, the Hit@1 of SubgraphRAG is 45.23, while it increases to 47.52 when using max pooling, resulting in an improvement of 2.23%. In contrast, using average pooling only yields a 1.9% increase.
@@ A.4 Prompts
The following is the detailed prompt template used in our reranking and reselection mechanism with path pooling for all experiments. Retrieved triple sequence has a large number of repeating entities. In order to form more coherent contextual and to save inference overhead, if, after reranking, the tail entity of the previous triple is the same as the head entity of the following triple, we merge the duplicate ones and using $\rightarrow$ to connect entities and relations.
Path Pooling Prompt System: Based on the paths retrieved from a knowledge graph, please answer the question. Please return formatted answers as a list, each prefixed with "ans:".   User: Paths:
m.011zsc4_->organization.leadership.organization->San Francisco Giants
m.0crtd80->sports.sports_league_participation.league->National League West …
San Francisco Giants->sports.sports_team.location->San Francisco
San Francisco Giants->time.participant.event->2014 Major League Baseball season
San Francisco Giants->time.participant.event->2012 Major League Baseball season
San Francisco Giants->sports.professional_sports_team.owner_s->Bill Neukom
San Francisco Giants->time.participant.event->2010 Major League Baseball season
San Francisco Giants->time.participant.event->2012 World Series
San Francisco Giants->time.participant.event->2010 World Series
San Francisco Giants->sports.sports_team.team_mascot->Crazy Crab
Lou Seal->sports.mascot.team->San Francisco Giants->sports.sports_team.championships->2014 World Series
Question: What year did the team with mascot named Lou Seal win the World Series?   Assistant: To find the year the team with mascot named Lou Seal won the World Series, we need to find the team with mascot named Lou Seal and then find the year they won the World Series. From the paths, we can see that Lou Seal is the mascot of the San Francisco Giants. Now, we need to find the year the San Francisco Giants won the World Series. From the paths, we can see that San Francisco Giants won the 2010 World Series and 2012 World Series and 2014 World Series. So, the team with mascot named Lou Seal (San Francisco Giants) won the World Series in 2010, 2012, and 2014. Therefore, the formatted answers are: ans: 2014 World Series ans: 2012 World Series ans: 2010 World Series   User: Paths: Question:

### v2

@@ A.3.1 Different Path Search Algorithms
Figure 11 summarizes Hit@1 and F1-Score of positional reranking mechanism with different path search algorithms on Qwen2.5-7b. Consistent with previous results, all three path search algorithms are effective in enhancing the performance of SubgraphRAG. As the number of triples increases, the Hit@1 and F1-Score of SubgraphRAG show a initially rising and then declining trend. However, when enhanced with a positional reranking mechanism based on BFS path search algorithm, both metrics demonstrate a consistently upward trend. For instance, with 500 triples, the Hit@1 of SubgraphRAG is 45.77, while it is 47.52 with BFS, which brings about an increase of 1.75%. The result emphasizes the effectiveness of incorporating the structure information of KGs.
@@ A.3.2 Different Pooling Strategies
As illustrated in Figure 12, positional reranking mechanism with different pooling strategies all enhanced SubgraphRAG on Qwen2.5-7b. Similar benefits are observed when leveraging triple reselection mechanism and on Llama3.1-8b. As commonly used pooling strategies, both max pooling and average pooling demonstrate significant roles here. However, surprisingly, when the number of triples increases from 50 to 200, the positional reranking mechanism using max pooling achieves better enhancement effect than using average pooling. For example, with 50 triples, the Hit@1 of SubgraphRAG is 45.23, while it increases to 47.52 when using max pooling, resulting in an improvement of 2.23%. In contrast, using average pooling only yields a 1.9% increase.